OpenStudy (anonymous):

(WILL Fan AND Medal!!) How to write a balanced net ionic equation for this equation.... h3po4(aq)+ca(oh)2(aq) --> ca3(po4)2(aq) + h2o(l) Appreciate your help!

OpenStudy (anonymous):

@abb0t do you understand this?

OpenStudy (abb0t):

STEPS: 1) Start by simply writing the overall balanced chemical reaction. 2) Then, you break apart the soluble molecules into the two ions that are formed (one positive and one negative). NOTE: Some things cannot be separated and you will have to use the solubility rules found in a textbook or found online to know. Or you can use reasoning. 3) Then, you simplify by canceling things out if they appear on both sides of the reaction, resulting in the Net Ionic Equation.

OpenStudy (anonymous):

Sooo... 2H3PO4 + 3Ca(OH)2 = Ca3(PO4)2 + 6H2O? (did it on paper)

OpenStudy (abb0t):

Good, now proceed to step 2 :)

OpenStudy (anonymous):

Oh, that was my conclusion :P is that not all the way balanced?

OpenStudy (abb0t):

Oh, I didn't do it. Lol. Um...give me a min. haha

OpenStudy (anonymous):

It's fine, take your time (:

OpenStudy (abb0t):

wait, is that what ur saying is ur net ionic?! or did u just balance it?

OpenStudy (anonymous):

lol it's just a balanced equation! Not the net ionic. haha

OpenStudy (abb0t):

Oh. I think that's right. loL. Well, your question asks for "net ionic". So I just assumed.

OpenStudy (anonymous):

That's what I'm wondering, the net ionic. :P Sorry, I was just stating before that I knew what the balanced equation is, just not the net ionic.

OpenStudy (abb0t):

Oh, net ionic meas you have 2 separate them by charge. for examle, 2CaCl2 means it's 2Ca 2+ and 2 Cl-

OpenStudy (anonymous):

So it would look like this? 2H3 + PO4 + 3Ca(OH)2 = Ca3(PO4)2 + 6H2O?

OpenStudy (anonymous):

I'm confused on the whole concept of "net ionic" and how to form one from a balanced equation. :S

OpenStudy (abb0t):

You have to look at the solubility rules, somethings don't separate. I can't remember them at the top of my head, but here is a website that will help: http://www.ausetute.com.au/solrules.html

OpenStudy (anonymous):

So if it's soluble, it separates, and if it isn't... it doesn't? Sorry if I'm talking gibberish here, I'm just trying to understand.